OpenStudy (anonymous):

What is the solution of the system? Use elimination. 3x-4y=9-3y+2y=9

OpenStudy (anonymous):

y=0 and x= 3 is the solution

OpenStudy (anonymous):

3x-4y=9 -3x+2y=9 Add them together

OpenStudy (anonymous):

x=3 and y=0 .

OpenStudy (anonymous):

this is not the original ones,if you sum you dont get y=0

OpenStudy (anonymous):

x=y=-9

OpenStudy (anonymous):

y=0 x=3 it is an easy question you can seperate the equations that way you will solve them easier